# ALINE: Joint Amortization for Bayesian Inference and Active Data Acquisition

### v1

@@ B.1 Training algorithm
20:              $\mathcal{L}=\mathcal{L}_{\text{NLL}}$ (Eq. 12) $+\mathcal{L}_{\text{PG}}$ (Eq. 11)
@@ B.2 Architecture and training details
The core of our architecture is a transformer network. We employ a configuration with 3 transformer layers, each equipped with 4 attention heads. The feedforward networks within each transformer layer have a dimension of 128. The model’s internal embedding dimension, consistent across the transformer layers and the output of the initial embedding layers, is 32. These transformer layers process the embedded representations of the context, query, and target sets. The interactions between these sets are governed by specific attention masks, visually detailed in Figure A2, where a shaded element indicates that the token corresponding to its row is permitted to attend to the token corresponding to its column.
Aline has two specialized output heads. The inference head, responsible for approximating posteriors and posterior predictive distributions, parameterizes a Gaussian Mixture Model (GMM) with 10 components. The embeddings corresponding to the inference targets are processed by 10 separate MLPs, one for each GMM component. Each MLP outputs parameters for its component: a mixture weight, a mean, and a standard deviation. The standard deviations are passed through a Softplus activation function to ensure positivity, and the mixture weights are normalized using a Softmax function. The policy head, which generates a probability distribution over the candidate query points, is a 2-layer MLP with a feedforward dimension of 128. Its output is passed through a Softmax function to ensure that the probabilities of all actions sum to unity. The architecture of Aline is shown in Figure A1.
@@
Figure A1: The Aline architecture. The model takes historical observations (Context Set), candidate query points (Query Set), and the current inference goal (Target Set) as inputs. These are transformed by Embedding Layers and subsequently by Transformer Layers. Finally, an Acquisition Head determines the next data point to query, while an Inference Head performs the approximate Bayesian inference.
@@ Appendix C Experimental details
This section provides details for the experimental setups. Section C.1 outlines the specifics for the active learning experiments in Section 5.1, including the synthetic function sampling procedures (Section C.1.1), implementation details for baseline methods (Section C.1.2), and training and evaluation details for these tasks (Section C.1.3). Next, in Section C.2 we describe the details of BED tasks, including the task descriptions (Section C.2.1), implementation of the baselines (Section C.2.2), and the training and evaluation details (Section C.2.3). Lastly, Section C.3 contains the specifics of the psychometric modeling experiments, detailing the psychometric function we use (Section C.3.1) and the setup for the experimental comparisons (Section C.3.2).
@@ C.1.1 Synthetic functions sampling procedure
The kernel’s output scale is sampled uniformly from the interval $U(0.1,1)$. The lengthscale(s) are sampled from $U(0.1,2)\times\sqrt{d_{x}}$. Input data points $x$ are sampled uniformly within the range $[-5,5]$ for each dimension. Finally, Gaussian noise with a fixed standard deviation of $0.01$ is added to the true function output $y$ for each sampled data point. Figure A3 illustrates some examples of the synthetic GP functions generated using this procedure.
@@ C.2.1 Task descriptions
Constant Elasticity of Substitution (CES) [1] considers a behavioral economics problem in which a participant compares two baskets of goods and rates the subjective difference in utility between the baskets on a sliding scale from 0 to 1. The utility of a basket $z$, consisting of $K$ goods with different values, is characterized by latent parameters $\theta=(\rho,\bm{\alpha},\mu)$. The design problem is to select pairs of baskets, $x=(z,z^{\prime})\in[0,100]^{2K}$, to infer the participant’s latent utility parameters.
@@ C.3.1 Model description
  • $\theta_{1}$ (threshold): The stimulus intensity at which the probability of a positive response reaches a specific criterion. It represents the location of the psychometric curve. We use a uniform prior $U[-3,3]$ for $\theta_{1}$.
@@ C.3.2 Experimental details
For each simulated experiment, true underlying parameters are sampled from their prior distributions. Stimulus values $x$ are selected from a discrete set of size 200 drawn uniformly from the range $[-5,5]$.
@@ Appendix D Additional experimental results
D.1 Active learning for regression and hyperparameter inference
@@
To further assess Aline, we present performance evaluations on an additional set of active learning benchmark functions, see Figure A4. The results on Gramacy and Branin show that we are on par with the GP baselines. For Three Hump Camel, we see both Aline and ACE-US showing reduced accuracy. This is because the function’s output value range extends beyond that of the GP functions used during pre-training. This highlights a potential area for future work, such as training Aline on a broader prior distribution of functions, potentially leading to more universally capable models.
@@
To qualitatively understand the behavior of our model, we visualize the query strategy employed by Aline for AL on a randomly sampled synthetic function Figure A5. This visualization illustrates how Aline iteratively selects query points to reduce uncertainty and refine its predictive posterior.
@@
We now visualize the evolution of Aline’s estimated posterior distributions for the underlying GP hyperparameters for a randomly drawn 2D synthetic GP function (see Figure A6). The posteriors are shown after 1, 15, and 30 active data acquisition steps. As Aline strategically queries more informative data points, its posterior beliefs about these generative parameters become increasingly concentrated and accurate.
@@
To assess the computational efficiency of Aline, we report the inference times for the AL tasks in Table A3. The times represent the total duration to complete a sequence of 30 steps for 1D functions and 50 steps for 2D functions, averaged over 10 independent runs. As both Aline and ACE-US perform inference via a single forward pass per step once trained, they are significantly faster compared to traditional GP-based methods.
@@
GP-US 0.62$\pm$0.09 1.72$\pm$0.23
GP-VR 1.41$\pm$0.14 4.03$\pm$0.18
GP-EPIG 1.34$\pm$0.11 3.43$\pm$0.24
ACE-US 0.08$\pm$0.00 0.19$\pm$0.02
Aline 0.08$\pm$0.00 0.19$\pm$0.02
D.2 Benchmarking on Bayesian experimental design tasks
In this section, we provide additional qualitative results for Aline’s performance on the BED benchmark tasks. Specifically, for the Location Finding task, we visualize the sequence of designs chosen by Aline and the resulting posterior distribution over the hidden source’s location (Figure A7). For the CES task, we present the estimated marginal posterior distributions for the model parameters, comparing them against their true underlying values (Figure A8). We see that Aline offers accurate parameter inference.
@@ D.2 Benchmarking on Bayesian experimental design tasks
D.3 Psychometric model
Targeted vs. full parameter acquisition.
To further demonstrate the advantage of Aline’s flexible targeting capability, we conduct an additional experiment within the psychometric modeling task. We evaluate two configurations: one where Aline specifically targets a subset of parameters (targeted), which we show in the main text, and another configuration where Aline aims to infer all four parameters simultaneously (full). As illustrated in Figure A9, directing Aline’s data acquisition strategy towards a specific parameter subset results in more accurate posterior inference for those targeted parameters compared to when Aline is tasked with learning all parameters concurrently. This underscores the benefit of targeted querying for achieving enhanced precision on parameters of primary interest.
Figure A9: Performance comparison for two Aline configurations under psychometric model. Aline (full) targets all four psychometric parameters simultaneously, and Aline (targeted) focuses on specific parameter subsets. When specifically targeting the guess and lapse rates, Aline (targeted) achieves significantly lower error for this subset compared to the Aline (full) version, demonstrating the benefit of its flexible focusing capability for improved parameter inference.
@@ Appendix E Computational resources and software
All experiments presented in this work, encompassing model development, hyperparameter optimization, baseline evaluations, and preliminary analyses, are performed on a GPU cluster equipped with AMD MI250X GPUs. The total computational resources consumed for this research, including all development stages and experimental runs, are estimated to be approximately 5000 GPU hours. For each experiment, it takes around 20 hours to train an Aline model for $10^{5}$ epochs. The core code base is built using Pytorch (https://pytorch.org/, License: modified BSD license). For the Gaussian Process (GP) based baselines, we utilize Scikit-learn [53] (https://scikit-learn.org/, License: modified BSD license). The DAD baseline is adapted from the original authors’ publicly available code [21] (https://github.com/ae-foster/dad; MIT License). Our implementation of the RL-BOED baseline uses the repository provided by [4] (https://github.com/yasirbarlas/RL-BOED; MIT License). We use questplus package (https://github.com/hoechenberger/questplus, License: GPL-3.0) to implement QUEST+, and use Psi-staircase (https://github.com/NNiehof/Psi-staircase, License: GPL-3.0) to implement the Psi-marginal method.

### v2

@@ B.1 Training algorithm
20:     $\mathcal{L}=\mathcal{L}_{\text{NLL}}$ (Eq. 12) $\mathcal{L}_{\text{PG}}$ (Eq. 11)
@@ B.2 Architecture and training details
The core of our architecture is a transformer network. We employ a configuration with 3 transformer layers, each equipped with 4 attention heads. The feedforward networks within each transformer layer have a dimension of 128. The model’s internal embedding dimension, consistent across the transformer layers and the output of the initial embedding layers, is 32. These transformer layers process the embedded representations of the context, query, and target sets. The interactions between these sets are governed by specific attention masks, visually detailed in Figure˜A1, where a shaded element indicates that the token corresponding to its row is permitted to attend to the token corresponding to its column.
Aline has two specialized output heads. The inference head, responsible for approximating posteriors and posterior predictive distributions, parameterizes a Gaussian Mixture Model (GMM) with 10 components. The embeddings corresponding to the inference targets are processed by 10 separate MLPs, one for each GMM component. Each MLP outputs parameters for its component: a mixture weight, a mean, and a standard deviation. The standard deviations are passed through a Softplus activation function to ensure positivity, and the mixture weights are normalized using a Softmax function. The policy head, which generates a probability distribution over the candidate query points, is a 2-layer MLP with a feedforward dimension of 128. Its output is passed through a Softmax function to ensure that the probabilities of all actions sum to unity. The architecture of Aline is shown in Figure˜2.
@@ Appendix C Experimental details
This section provides details for the experimental setups. Section˜C.1 outlines the specifics for the active learning experiments in Section˜4.1, including the synthetic function sampling procedures (Section˜C.1.1), implementation details for baseline methods (Section˜C.1.2), and training and evaluation details for these tasks (Section˜C.1.3). Next, in Section˜C.2 we describe the details of BED tasks, including the task descriptions (Section˜C.2.1), implementation of the baselines (Section˜C.2.2), and the training and evaluation details (Section˜C.2.3). Lastly, Section˜C.3 contains the specifics of the psychometric modeling experiments, detailing the psychometric function we use (Section˜C.3.1) and the setup for the experimental comparisons (Section˜C.3.2).
@@ C.1.1 Synthetic functions sampling procedure
The kernel’s output scale is sampled uniformly from the interval $U(0.1,1)$. The lengthscale(s) are sampled from $U(0.1,2)\times\sqrt{d_{x}}$. Input data points $x$ are sampled uniformly within the range [5,5][-5,5] for each dimension. Finally, Gaussian noise with a fixed standard deviation of $0.01$ is added to the true function output $y$ for each sampled data point. Figure˜A2 illustrates some examples of the synthetic GP functions generated using this procedure.
@@ C.2.1 Task descriptions
Constant Elasticity of Substitution (CES) [3] considers a behavioral economics problem in which a participant compares two baskets of goods and rates the subjective difference in utility between the baskets on a sliding scale from 0 to 1. The utility of a basket $z$, consisting of $K$ goods with different values, is characterized by latent parameters $\theta=(\rho,\bm{\alpha},u)$. The design problem is to select pairs of baskets, $x=(z,z^{\prime})\in[0,100]^{2K}$, to infer the participant’s latent utility parameters.
@@ C.2.2 Implementation details of baselines
vsOED [65] is an amortized BED method that employs an actor-critic reinforcement learning framework. It utilizes separate networks for the actor (policy), the critic (value function), and the variational posterior approximation. The reward signal for training the policy is the incremental improvement in the log-probability of the ground-truth parameters under the variational posterior, which is estimated at each step of the experiment. Following the original implementation, a distinct posterior network is trained for each design stage, while the actor and critic share a common backbone. For our implementation, the hidden layers of all networks are 3-layer MLPs with 256 units and ReLU activations. The posterior network outputs the parameters for an 8-component Gaussian Mixture Model (GMM). The input to the actor and critic networks is the zero-padded history of design-observation pairs, concatenated with a one-hot encoding of the current time step. We train for $10^{4}$ epochs with a batch size of $10^{4}$ and a $10^{6}$-sized replay buffer. The learning rate starts at $10^{-3}$ with a 0.9999 exponential decay per epoch, and the discount factor is 0.9. To encourage exploration for the deterministic policy, Gaussian noise is added during training; the initial noise scale is 0.5 for the Location Finding task and 5.0 for the CES task, with decay rates of 0.9999 and 0.9998, respectively.
@@ C.3.1 Model description
  • $\theta_{1}$ (threshold): The stimulus intensity at which the probability of a positive response reaches a specific criterion. It represents the location of the psychometric curve. We use a uniform prior U[3,3]U[-3,3] for $\theta_{1}$.
@@ C.3.2 Experimental details
For each simulated experiment, true underlying parameters are sampled from their prior distributions. Stimulus values $x$ are selected from a discrete set of size 200 drawn uniformly from the range [5,5][-5,5].
@@ Appendix D Additional experimental results
D.1 Active Exploration of High-Dimensional Hyperparameter Landscapes
To demonstrate Aline’s utility on complex, high-dimensional tasks, we conduct a new set of experiments on actively exploring hyperparameter performance landscapes. This experiment aims to efficiently characterize a machine learning model’s overall behavior on a new task, allowing practitioners to quickly assess a model family’s viability or understand its sensitivities. The task is to actively query a small number of hyperparameter configurations to build a surrogate model that accurately predicts performance for a larger, held-out set of target configurations. We use high-dimensional, real-world tasks from the HPO-B benchmark [2], evaluating on rpart (6D), svm (8D), ranger (9D), and xgboost (16D) datasets. Aline is trained on their multiple pre-defined training sets. We then evaluate its performance, alongside non-amortized GP-based baselines and an amortized surrogate baseline (ACE-US), on the benchmark’s held-out and entirely unseen test sets.
Table A3 shows the RMSE results after 30 steps, averaged across all test tasks for each dataset. First, both amortized methods, Aline and ACE-US, significantly outperform all non-amortized GP-based baselines across all tasks. This highlights the power of meta-learning in this domain. GP-based methods must learn each new performance landscape from scratch, which is highly inefficient in high dimensions. In contrast, both Aline and ACE-US are pre-trained on hundreds of related tasks, and their Transformer architectures meta-learn the structural patterns common to these landscapes. This shared prior knowledge allows them to make far more accurate predictions from sparse data. Second, while ACE-US performs strongly due to its amortized nature, Aline consistently achieves the best or joint-best performance. This demonstrates the additional, crucial benefit of our core contribution: the learned acquisition policy. ACE-US relies on a standard heuristic, whereas Aline’s policy is trained end-to-end to learn how to optimally explore the landscape, leading to more informative queries and ultimately a more accurate final surrogate model.
Table A3: RMSE ($\downarrow$) on the HPO-B benchmark after 30 active queries. Results show the mean and 95% CI over all test tasks for each dataset.
GP-RS GP-US GP-VR GP-EPIG ACE-US Aline (ours)
rpart (6D) $07\pm 0.03 0.0$$4\pm 0.02 0.0$$4\pm 0.02 0.0$$5\pm 0.02 0.0$$\mathbf{0.01}\pm 0$.00 $\mathbf{0.01\pm}0.00$
[…]
D.2 Active learning for regression and hyperparameter inference
@@
To further assess Aline, we present performance evaluations on an additional set of active learning benchmark functions, see Figure˜A3. The results on Gramacy and Branin show that we are on par with the GP baselines. For the Three Hump Camel, we see both Aline and ACE-US showing reduced accuracy. This is because the function’s output value range extends beyond that of the GP functions used during pre-training. This highlights a potential area for future work, such as training Aline on a broader prior distribution of functions, potentially leading to more universally capable models.
@@
To qualitatively understand the behavior of our model, we visualize the query strategy employed by Aline for AL on a randomly sampled synthetic function Figure˜A4. This visualization illustrates how Aline iteratively selects query points to reduce uncertainty and refine its predictive posterior.
@@
We now visualize the evolution of Aline’s estimated posterior distributions for the underlying GP hyperparameters for a randomly drawn 2D synthetic GP function (see Figure˜A5). The posteriors are shown after 1, 15, and 30 active data acquisition steps. As Aline strategically queries more informative data points, its posterior beliefs about these generative parameters become increasingly concentrated and accurate.
@@
To assess the computational efficiency of Aline, we report the inference times for the AL tasks in Table˜A4. The times represent the total duration to complete a sequence of 30 steps for 1D functions and 50 steps for 2D functions, averaged over 10 independent runs. As both Aline and ACE-US perform inference via a single forward pass per step once trained, they are significantly faster compared to traditional GP-based methods.
@@
[…]
D.3 Benchmarking on Bayesian experimental design tasks
In this section, we provide additional qualitative results for Aline’s performance on the BED benchmark tasks. Specifically, for the Location Finding task, we visualize the sequence of designs chosen by Aline and the resulting posterior distribution over the hidden source’s location (Figure˜A6). For the CES task, we present the estimated marginal posterior distributions for the model parameters, comparing them against their true underlying values (Figure˜A7). We see that Aline offers accurate parameter inference.
@@ D.3 Benchmarking on Bayesian experimental design tasks
D.4 Psychometric model
Demonstrations of flexibility.
We conduct two ablations to explicitly validate Aline’s flexible targeting capabilities.
First, we test the ability to switch targets mid-rollout. We configure a single experiment where for the first 15 steps, the target is threshold & slope parameters, and at step 16, the target is switched to the guess rate & lapse rate. As shown in Figure A8(a), Aline’s acquisition strategy adapts immediately and correctly, shifting its queries from the decision threshold region to the extremes of the stimulus range to gain maximal information about the new targets.
Second, we test generalization to novel target combinations. A single Aline model is trained to handle two distinct targets separately: (1) threshold & slope and (2) guess & lapse rate. At deployment, we task this model with a novel, unseen combination: targeting all four parameters simultaneously. As shown in Figure A8(b), the resulting policy is a sensible mixture of the two underlying strategies it has learned, strategically alternating queries between points near the decision threshold and points at the extremes. This confirms that Aline can successfully compose its learned strategies to generalize to new inference goals at runtime.
Figure A8: Demonstration of Aline’s runtime flexibility on the psychometric task. (a) The acquisition strategy adapts after the inference target is switched mid-rollout from (threshold & slope) to (guess & lapse rate). (b) When tasked with a novel combined target (all four parameters), the policy generalizes by mixing the two distinct strategies it learned during training.
@@ Appendix E Computational resources and software
All experiments presented in this work, encompassing model development, hyperparameter optimization, baseline evaluations, and preliminary analyses, are performed on a GPU cluster equipped with AMD MI250X GPUs. The total computational resources consumed for this research, including all development stages and experimental runs, are estimated to be approximately 5000 GPU hours. For each experiment, it takes around 20 hours to train an Aline model for $10^{5}$ epochs. The core code base is built using Pytorch (https://pytorch.org/, License: modified BSD license). For the Gaussian Process (GP) based baselines, we utilize Scikit-learn [56] (https://scikit-learn.org/, License: modified BSD license). The DAD baseline is adapted from the original authors’ publicly available code [23] (https://github.com/ae-foster/dad; MIT License). Our implementations of the RL-BOED and vsOED baselines are adapted from the official repositories provided by [6] (https://github.com/yasirbarlas/RL-BOED; MIT License) and [65] (https://github.com/wgshen/vsOED; MIT License), respectively. We use questplus package (https://github.com/hoechenberger/questplus, License: GPL-3.0) to implement QUEST+, and use Psi-staircase (https://github.com/NNiehof/Psi-staircase, License: GPL-3.0) to implement the Psi-marginal method.